# Inflationary Flows: Calibrated Bayesian Inference with Diffusion-Based Models

### v1

@@ 1 Introduction
Alternatively, methods based on variational inference (VI) attempt to approximate posterior distributions by optimization, minimizing some measure of divergence between the true posterior and a parameterized set of distributions $q_{\phi}(\mathbf{z}|\mathbf{x})$ [2]. For example, methods like the variational autoencoder (VAE) [3, 4] minimize the Kullback-Leibler (KL) divergence between true and approximate posteriors, producing bidirectional mappings between data and latent spaces. However, despite their many successes, these models tend to produce blurry data reconstructions and non-unique latent spaces without additional assumptions [5, 6, 7]. Similarly, posterior uncertainty estimates are overconfident due to minimization of the reverse (mode-seeking) KL divergence [8, 9]. By contrast, normalizing flow (NF) models [10, 11] work by applying a series of bijective transformations to a simple base distribution (usually uniform or Gaussian) to deterministically convert samples to a desired target distribution. While NFs have been successfully used for posterior approximation [12, 13, 14, 15, 16] and produce higher-quality samples, the requirement that the Jacobian of each transformation be simple to compute often requires a high number of transformations, and these transformations do not alter the the dimensionality of their inputs, resulting in latent spaces with thousands of dimensions. Thus, these models lack a key advantage of VAEs and their ilk: a compressed representation of the data.
@@
More recently, diffusion-based models (DBMs) [17, 18, 19, 20, 21, 22, 23, 24] have been shown to achieve state-of-the-art results in several generative tasks, including image, sound, and text-to-image generation. These models work by stipulating a fixed forward noising process (i.e., forward SDE), wherein Gaussian noise is incrementally added to samples of the target data distribution until all information in the original data is degraded. To generate samples from the target distribution, one then needs to simulate the reverse de-noising process (reverse SDE [25]) which requires knowledge of the score of the intermediate “noised” transitional densities. Estimation of this score function across multiple noise levels is the key component of DBM model training, typically using a de-noising score matching objective [26, 19, 21]. Yet, despite their excellent performance as generative models, DBMs, unlike VAEs or NFs, do not readily lend themselves to inference. In particular, because DBMs use a diffusion process to transform the data distribution, they fail to preserve local structure in the data (Figure 1), and uncertainty under this mapping is high at its endpoint because of continuous noise injection and resultant mixing. Moreover, because the final distribution—Gaussian white noise of the same dimension—must have higher entropy than the original data, there is no data compression.
@@
Thus, despite tremendous improvements in sample quality, modern generative models do not lend themselves to one of the key modeling goals in scientific applications: accurate Bayesian inference. Here, we address this challenge by demonstrating how a novel DBM variant that we call inflationary flows can, in fact, produce calibrated, identifiable Bayesian inference. Specifically, our contributions are: First, we show how a previously established link between the SDE defining diffusion models and the probability flow ODE (pfODE) that gives rise to the same Fokker-Planck equation [21] can be used to define a unique, deterministic map between the original data and an asymptotically Gaussian distribution. This map is bidirectional, preserves local neighborhoods, and has controllable numerical error, making it suitable for rigorous uncertainty quantification. Second, we define two classes of flows that correspond to novel noise injection schedules in the forward SDE of the diffusion model. The first of these preserves a measure of dimensionality, the participation ratio (PR) [27], based on second-order data statistics, preventing an effective increase in data dimensionality with added noise, while the second flow reduces PR, providing data compression. Finally, we demonstrate experimentally that inflationary flows indeed preserve local neighborhood structure, allowing for sampling-based uncertainty estimation, and that these models continue to provide high-quality generation under compression, even from latent spaces reduced to as little as 2% of the nominal data dimensionality. As a result, inflationary flows offer excellent generative performance while affording data compression and accurate uncertainty estimation for scientific applications.
@@ 2 Three views of diffusion-based models
Figure 1: SDE-ODE Duality of diffusion-based models. The forward (noising) SDE defining the DBM (left) gives rise to a sequence of marginal probability densities whose temporal evolution is described by a Fokker-Planck equation (FPE, middle). But this correspondence is not unique: the probability flow ODE (pfODE, right) gives rise to the same FPE. That is, while both the SDE and the pfODE transform the data distribution in the same way, the former is noisy and mixing while the latter is deterministic and neighborhood-preserving. Both models require knowledge of the score function $\nabla_{\mathbf{x}}\log p_{t}(\mathbf{x})$, which can learned by training either.
As with standard DBMs, we assume a data distribution $p_{0}(\mathbf{x})=p_{\mathrm{data}}(\mathbf{x})$ at time $t=0$, transformed via a forward noising process defined by the stochastic differential equation [e.g., 17, 19]:
@@
Unlike (1), this process is deterministic, and data points evolve smoothly (Figure 1, right), resulting in a flow that preserves local neighborhoods. Moreover, the pfODE is uniquely defined by $\mathbf{f}(\mathbf{x},t)$, $\mathbf{G}(\mathbf{x},t)$, and the score function. This connection between the marginals satisfying the SDEs of diffusion processes and deterministic flows described by an equivalent ODE has also been recently explored in the context of flow matching models [30, 31, 32, 33, 34, 35], a connection on which we elaborate in Section 7.
@@ 3 Inflationary flows
It is straightforward to show (Appendix A.1) that this class of time-varying densities satisfies (3) when $\mathbf{f}=\mathbf{0}$ and $\mathbf{GG^{\top}}=\mathbf{\dot{C}}$. This can be viewed as a process of deterministically “inflating” each point in the data set, or equivalently as smoothing the underlying data distribution on ever coarser scales, similar to denoising approaches to DBMs [36, 37]. Eventually, if the smoothing kernel grows much larger than $\boldsymbol{\Sigma}_{0}$, the covariance in the original data, total covariance $\boldsymbol{\Sigma}(t)\equiv\boldsymbol{\Sigma}_{0}+\mathbf{C}(t)\rightarrow\mathbf{C}(t)$, $p_{t}(\mathbf{x})\approx\mathcal{N}(\mathbf{0},\mathbf{C}(t))$, and all information has been removed from the original distribution. However, because it is numerically inconvenient for the variance of the asymptotic distribution $p_{\infty}(\mathbf{x})$ to grow much larger than that of the data, we follow previous work in adding a time-dependent coordinate rescaling $\mathbf{\tilde{x}}(t)=\mathbf{A}(t)\cdot\mathbf{x}(t)$ [21, 28], resulting in an asymptotic solution $p_{\infty}(\mathbf{x})=\mathcal{N}(\mathbf{0},\mathbf{A}\boldsymbol{\Sigma}\mathbf{A^{\top}})$ of the corresponding Fokker-Planck equation when $\boldsymbol{\dot{\Sigma}}=\mathbf{\dot{C}}$ and $\mathbf{\dot{A}}\boldsymbol{\Sigma}\mathbf{A}^{\top}+\mathbf{A}\boldsymbol{\Sigma}\mathbf{\dot{A}}^{\top}=\mathbf{0}$ (Appendix A.2). Together, these assumptions give rise to the pfODE (Appendix A.3):
@@ 3.2 Dimension-reducing flows
In the previous section, we saw that isotropic inflation preserves intrinsic data dimensionality as measured by PR. Here, we generalize and consider anisotropic inflation at different rates along each of the eigenvectors of $\boldsymbol{\Sigma}$: $\tfrac{\mathrm{d}}{\mathrm{d}t}(\boldsymbol{{\sigma}^{2}})=\rho\mathbf{g}\odot\boldsymbol{\sigma^{2}}$ with $g_{*}\equiv\max(\mathbf{g})$ so that the fastest inflation rate is $\rho g_{*}$. Then, if we take $g_{i}=g_{*}$ for $i\in\{i_{1},i_{2},\ldots i_{K}\}$ and $g_{i}<g_{*}$ for the other dimensions,
@@
which is the dimension that would be achieved by simply truncating the original covariance matrix in a manner set by our choice of $\mathbf{g}$. Here, unlike in (9), we do not aim for rescaling to compensate for expansion along each dimension, since that would undo the effect of differential inflation rates. Instead, we choose a single global rescaling factor $\alpha(t)\propto A_{0}\exp(-\rho g_{*}t/2)$, which converges to a Gaussian distribution with covariance matching that of the initial data in all dimensions with $g_{i}=g_{*}$.
@@
Figure 3: Dimension-reducing flows for toy datasets. Numerical simulations of dimension-reducing flows for the same five datasets as in Figure 2. For 2D datasets, we showcase reduction from two to one dimension, while 3D datasets are reduced to two dimensions. Colors and layouts are the same as in Figure 2, with scores again estimated using neural networks trained on each example. Additional results showcasing (1) similar flows further compressing two-dimensional manifolds embedded in $D=3$ space, and (2) effects of adopting different scaling schemes for target data are given in Appendices C.2.1 and C.2.2, respectively.
@@ 4 Score function approximation from DBMs
Briefly, in line with previous work on DBMs [28], we train neural networks to estimate a de-noised version $\mathbf{D}(\mathbf{x},\mathbf{C}(t))$ of a noise-corrupted data sample $\mathbf{x}$ given noise level $\mathbf{C}(t)$ (cf. Appendix A.4 for the correspondence between $\mathbf{C}(t)$ and noise). That is, we model $\mathbf{D}_{\theta}(\mathbf{x},\mathbf{C}(t))$ using a neural network and train it by minimizing a standard $L_{2}$ de-noising error:
@@ 5 Calibrated uncertainty estimates from inflationary flows
The only potential source of error, apart from Monte Carlo error, in the above procedure arises from the fact that the score function used in (6) is only an estimate of the true score. While it is possible that integrating noisy estimates of the score could produce errant sample mappings, results shown in Figures 4B–C suggest that such numerical errors do not seem to accumulate in cases where they can be compared with exact results. This is because, empirically, score estimates do not appear to be strongly auto-correlated in time (Appendix C.3), suggesting that $\mathbf{\hat{s}}(\mathbf{x},t)$, the score estimate, is well approximated as a scaled colored noise process and the corresponding pfODE as an SDE. In such a case, standard theorems for SDE integration show that while errors due to noise do accumulate, these can be mitigated by a careful choice of integrator and ultimately controlled by reducing step size [41, 42]. Below, we verify this empirically in both low-dimensional examples and with round-trip integration of the pfODE in high-dimensional datasets.
@@ 6 Experiments
Figure 4: Calibration experiments. For select toy datasets, we numerically assessed coverage during the inflation and generation procedures (Appendix B.6). (A) We constructed fixed coverage sets by sampling data points at fixed Mahalanobis radii from the centers of each distribution and creating alpha shapes (2D) or meshes (3D). (B–C) We then quantified the change in coverage fraction for each of these sets at the end of either process. Lines represent means and shaded regions ±$\pm 2$ standard deviations across three sets of random seeds. (D) Illustration of the effect of flows on set geometry. While both types of flows distort the shapes of initial sets, they do preserve local neighborhoods, even when one dimension is compressed by five orders of magnitude.
@@
Figure 5: Generation and Round-Trip Experiments for CIFAR-10. Top row: Generated samples for each flow schedule (PR-Preserving (PRP), and PR-Reducing to 2%, 10%, and 20%). Bottom row: Results for round-trip experiments. Leftmost columns are original samples, middle columns are samples mapped to Gaussian latent spaces, and rightmost columns are recovered samples.
To verify that inflationary flows successfully compress high-dimensional data, we compressed two benchmark datasets (CIFAR-10 [43], AFHQv2 [44]) to 20%, 10%, and 2% of their nominal dimensionality and examined their round-trip and pure generative performance. For estimating score, we used DBM networks trained using the scheme of [28] combined with the novel scaling, noise, and preconditioning proposed above (see Appendices B.1, B.4.2 for details; code available at [45]).
Figure 6: Generation and Round-Trip Experiments for AFHQv2. Layout and setup same as for Figure 5 - see Appendices B.5.2, B.5.1 for details.
@@
Table 1: FID and Round-Trip MSE Mean ±$\pm 2\sigma$ for benchmark datasets
Metric Dataset PRP PRR to 2% PRR to 10% PRR to 20%
FID CIFAR-10 $17.01\pm 0.10$ $22.23\pm 0.16$ $23.63\pm 0.13$ $25.93\pm 0.40$
Round-Trip MSE CIFAR-10 $0.23\pm 0.03$ $2.06\pm 0.04$ $2.25\pm 0.01$ $4.16\pm 0.39$
FID AFHQv2 $11.89\pm 0.08$ $13.07\pm 0.07$ $13.67\pm 0.09$ $16.77\pm 0.14$
Round-Trip MSE AFHQv2 $0.38\pm 0.04$ $5.57\pm 0.20$ $7.95\pm 0.31$ $8.17\pm 0.08$
In Figures 5 and 6, we show 24 randomly generated target distribution samples for each schedule (PR-Preserving (PRP), PR-Reducing to 20%, 10%, and 2% (PRR to 20%, 10%, 2%)) and benchmark image dataset (CIFAR-10, AFHQv2). In addition, we show computed Frechet Inception Distance (FID) scores [46] over three different sets of 50,000 randomly generated images for each of these schedules and for both datasets (Table 1). Note that preserving intrinsic dimension (PRP schedules) leads to better (smaller) FID scores, as expected. Perhaps surprisingly, increasing the number of preserved dimensions in the PR-Reducing regimes leads to worse (i.e., higher) FID scores. This is because retaining more dimensions in our PR-Reducing schedules leads to larger scale gaps between our preserved and compressed dimensions (i.e., larger $e^{\rho(g_{*}-g_{i})T}$), thus increasing the required noise range over which networks must optimally estimate scores. That is, PR-Reducing schedules with higher numbers of preserved-dimensions pose a more challenging learning problem (Appendix B.2).
Additionally, Figures 5 and 6 also show results for round-trip integration (from data to compressed space and back) for eight random samples across the same experimental configurations and datasets. Reconstructions are qualitatively good across all conditions. Of note, these networks produce inflated representations that resemble true low-rank Gaussian samples, consistent with a reduced-dimensional latent space (see PR-Reducing to 2%, middle columns). In Table 1, we also include mean squared errors (MSEs) for each such round-trip experiment (across all schedules and datasets), computed over 3 randomly sampled sets, each with 10K images. Trends observed here are similar to the ones seen for FID experiments. PR-Preserving networks produced lower MSEs overall, whereas PR-Reducing networks with higher percentage of preserved dimensions (i.e., 20%) yielded higher MSEs than PR-Reducing networks with smaller number of preserved dimensions (i.e., 2%). Together, these results suggest that dimension reduction with inflationary flows may necessitate trade-offs between effective compression and the difficulty of score estimation, as noted above.

### v3

@@ 1 Introduction
Alternatively, methods based on variational inference (VI) attempt to approximate posterior distributions by optimization, minimizing some measure of divergence between the true posterior and a parameterized set of distributions $q_{\phi}(\mathbf{z}|\mathbf{x})$ [2]. For example, methods like the variational autoencoder (VAE) [3, 4] minimize the Kullback-Leibler (KL) divergence between true and approximate posteriors, producing bidirectional mappings between data and latent spaces. In vanilla VAEs, posterior uncertainty estimates are typically overconfident due to minimization of the reverse (mode-seeking) KL divergence [5, 6]. While some lines of work have sought to mitigate this posterior mismatch problem by utilizing different divergences 7, 8, 9, 10, VAEs still tend to produce blurry data reconstructions and non-unique latent spaces without additional assumptions [11, 12, 13].
By contrast, normalizing flow (NF) models [14, 15] work by applying a series of bijective transformations to a simple base distribution (usually uniform or Gaussian) to deterministically convert samples to a desired target distribution. While NFs have been successfully used for posterior approximation [16, 17, 18, 19, 20] and produce higher-quality samples, the requirement that the Jacobian of each transformation be simple to compute often requires a high number of transformations and, traditionally, these transformations do not alter the the dimensionality of their inputs, resulting in latent spaces with thousands of dimensions. More recent lines of work on injective flow models 21, 22, 23, 24, 25 address this limitation by allowing practitioners to use flows to learn lower dimensional manifolds from data, but most compression-capable flow models still fail to reach high generative performance on key benchmark image datasets (cf. [23]).
@@
More recently, diffusion-based models (DBMs) [26, 27, 28, 29, 30, 31, 32, 33] have been shown to achieve state-of-the-art results in several generative tasks, including image, sound, and text-to-image generation. These models work by stipulating a fixed forward noising process (e.g., a forward stochastic differential equation (SDE)), wherein Gaussian noise is incrementally added to samples of the target data distribution until all information in the original data is degraded. To generate samples from the target distribution, one then needs to simulate the reverse de-noising process (reverse SDE [34]) which requires knowledge of the score of the intermediate “noised” transitional densities. Estimation of this score function across multiple noise levels is the key component of DBM model training, typically using a de-noising score matching objective [35, 28, 30]. Yet, despite their excellent performance as generative models, DBMs, unlike VAEs or flows, do not readily lend themselves to inference. In particular, because DBMs use a diffusion process to transform the data distribution, they fail to preserve local structure in the data (Figure 1), and uncertainty under this mapping is high at its endpoint because of continuous noise injection and resultant mixing. Moreover, because the final distribution—Gaussian white noise of the same dimension—must have higher entropy than the original data, there is no data compression.
Finally, emerging work on flow matching models [36, 37, 38, 39, 40, 41, 42] has achieved impressive generative performance on several benchmark image datasets. Such models utilize simple conditional distribution families to learn a vector field capable of transporting points between two pre-specified densities. These are closely related to the probability flow ODE (pfODE) view of DBMs, and, in fact, have been shown to be equivalent to such models for specific choices of “interpolant” functions and conditional distributions. Despite their exceptional generative performance and deterministic nature, existing flow matching approaches do not allow for compression and, therefore, do not allow practitioners to infer a lower dimensional latent space from data.
Thus, despite tremendous improvements in sample quality, modern generative models do not lend themselves to one of the key modeling goals in scientific applications: calibrated Bayesian inference. Note that while many works focus on predictive calibration, how well the inferred marginal $p(\mathbf{x})$ matches real data [43, 44, 45, 46, 47], our focus here is on posterior calibration, how well $q(\mathbf{z}|\mathbf{x})$ matches the true posterior $p(\mathbf{z}|\mathbf{x})$. We address this challenge by demonstrating how a novel DBM variant that we call inflationary flows can, in fact, produce calibrated Bayesian inference in this sense.
@@
Specifically, our contributions are: First, focusing on the case of unconditional generative models, we show how a previously established link between the SDE defining diffusion models and the probability flow ODE (pfODE) that gives rise to the same Fokker-Planck equation [30] can be used to define a unique, deterministic map between the original data and an asymptotically Gaussian distribution. This map is bidirectional, preserves local neighborhoods, and has controllable numerical error, making it suitable for rigorous uncertainty quantification. Second, we define two classes of flows that correspond to novel noise injection schedules in the forward SDE of the diffusion model. The first of these preserves a measure of dimensionality, the participation ratio (PR) [48], based on second-order data statistics, preventing an effective increase in data dimensionality with added noise, while the second flow reduces PR, providing data compression. We demonstrate experimentally that inflationary flows indeed preserve local neighborhood structure, allowing for sampling-based uncertainty estimation, and that these models continue to provide high-quality generation under compression, even from latent spaces reduced to as little as 0.03% of the nominal data dimensionality. As a result, inflationary flows offer excellent generative performance while affording data compression and accurate uncertainty estimation for scientific applications.
@@ 2 Three views of diffusion-based models
Figure 1: SDE-ODE Duality of diffusion-based models. The forward (noising) SDE defining the DBM (left) gives rise to a sequence of marginal probability densities whose temporal evolution is described by a Fokker-Planck equation (FPE, middle). But this correspondence is not unique: the probability flow ODE (pfODE, right) gives rise to the same FPE. That is, while both the SDE and the pfODE possess the same marginals, the former is noisy and mixing while the latter is deterministic and neighborhood-preserving. Both models require knowledge of the score function $\nabla_{\mathbf{x}}\log p_{t}(\mathbf{x})$, which can learned by training either model.
As with standard DBMs, we assume a data distribution $p_{0}(\mathbf{x})=p_{\mathrm{data}}(\mathbf{x})$ at time $t=0$, transformed via a forward noising process defined by the stochastic differential equation [e.g., 26, 28]:
@@
Unlike (1), this process is deterministic, and data points evolve smoothly (Figure 1, right), resulting in a flow that preserves local neighborhoods. Moreover, the pfODE is uniquely defined by $\mathbf{f}(\mathbf{x},t)$, $\mathbf{G}(\mathbf{x},t)$, and the score function. This connection between the marginals satisfying the SDEs of diffusion processes and deterministic flows described by an equivalent ODE has also been recently explored in the context of flow matching models [36, 37, 38, 39, 40, 41, 42], a connection on which we elaborate in Section 7.
@@ 3 Inflationary flows
It is straightforward to show (Appendix A.1) that this class of time-varying densities satisfies (3) when $\mathbf{f}=\mathbf{0}$ and $\mathbf{GG^{\top}}=\mathbf{\dot{C}}$. This can be viewed as a process of deterministically “inflating” each point in the data set, or equivalently as smoothing the underlying data distribution on ever coarser scales, similar to denoising approaches to DBMs [51, 52]. Eventually, if the smoothing kernel grows much larger than $\boldsymbol{\Sigma}_{0}$, the covariance in the original data, total covariance $\boldsymbol{\Sigma}(t)\equiv\boldsymbol{\Sigma}_{0}+\mathbf{C}(t)\rightarrow\mathbf{C}(t)$, $p_{t}(\mathbf{x})\approx\mathcal{N}(\mathbf{0},\mathbf{C}(t))$, and all information has been removed from the original distribution. However, because it is numerically inconvenient for the variance of the asymptotic distribution $p_{\infty}(\mathbf{x})$ to grow much larger than that of the data, we follow previous work in adding a time-dependent coordinate rescaling $\mathbf{\tilde{x}}(t)=\mathbf{A}(t)\cdot\mathbf{x}(t)$ [30, 49], which results in an asymptotic solution $p_{\infty}(\mathbf{x})=\mathcal{N}(\mathbf{0},\mathbf{A}\boldsymbol{\Sigma}\mathbf{A^{\top}})$ of the corresponding Fokker-Planck equation when $\boldsymbol{\dot{\Sigma}}=\mathbf{\dot{C}}$ and $\mathbf{\dot{A}}\boldsymbol{\Sigma}\mathbf{A}^{\top}+\mathbf{A}\boldsymbol{\Sigma}\mathbf{\dot{A}}^{\top}=\mathbf{0}$ (Appendix A.2). Together, these assumptions give rise to the pfODE (Appendix A.3):
@@ 3.2 Dimension-reducing flows
In the previous section, we saw that isotropic inflation preserves intrinsic data dimensionality as measured by PR. Here, we generalize and consider anisotropic inflation at different rates along each of the eigenvectors of $\boldsymbol{\Sigma}$: $\tfrac{\mathrm{d}}{\mathrm{d}t}(\boldsymbol{{\sigma}^{2}})=\rho\mathbf{g}\odot\boldsymbol{\sigma^{2}}$. In addition, we denote $g_{*}\equiv\max(\mathbf{g})$, so that the fastest inflation rate is $\rho g_{*}$. Then, if we take $g_{i}=g_{*}$ for $i\in\{i_{1},i_{2},\ldots i_{K}\}$ and $g_{i}<g_{*}$ for the other dimensions,
@@
which is the dimension that would be achieved by simply truncating the original covariance matrix in a manner set by our choice of $\mathbf{g}$. Here, unlike in (9), we do not aim for rescaling to compensate for expansion along each dimension, since that would undo the effect of differential inflation rates. Instead, we choose a single global rescaling factor $\alpha(t)\propto A_{0}\exp(-\rho g_{*}t/2)$, leading to a Gaussian asymptotic solution with the original data covariance in dimensions $i\in\{i_{1},i_{2},\ldots i_{K}\}$.
@@
Figure 3: Dimension-reducing flows for toy datasets. Numerical simulations of dimension-reducing flows for the same five datasets as in Figure 2. For 2D datasets, we showcase reduction from two to one dimension, while 3D datasets are reduced to two dimensions. Colors and layouts are the same as in Figure 2, with scores again estimated using neural networks trained on each example. Additional results showcasing (1) similar flows further compressing two-dimensional manifolds embedded in $D=3$ space, and (2) effects of adopting different scaling schemes for target data are given in Appendices C.2.2 and C.2.3, respectively.
@@ 4 Score function approximation from DBMs
Briefly, in line with previous work on DBMs [49], we train neural networks to estimate a de-noised version, $\mathbf{D}(\mathbf{x},\mathbf{C}(t))$, of a noise-corrupted data sample $\mathbf{x}$ given noise level $\mathbf{C}(t)$ (cf. Appendix A.4 for the correspondence between $\mathbf{C}(t)$ and noise). That is, we model $\mathbf{D}_{\theta}(\mathbf{x},\mathbf{C}(t))$ using a neural network and train it by minimizing a standard $L_{2}$ de-noising error:
@@ 5 Calibrated uncertainty estimates from inflationary flows
Several previous lines of work [43, 44, 45, 46, 47] have focused on assessing how well model-predicted marginals $p(\mathbf{x})$ match real data (i.e., the predictive calibration case). Though we do compare our models’ predictive calibration performance against existing injective flow models (Table 3), here we are primarily focused on quantifying error in unconditional posterior inference. That is, we are interested in quantifying the mismatch between inferred posteriors $q(\mathbf{z}|\mathbf{x})$ and true posteriors $p(\mathbf{z}|\mathbf{x})$, especially in contexts where the true generative model is unknown and must be learned from data. This is by far the most common scenario in modern generative models like VAEs, flows, and GANs.
@@
Figure 4: Calibration experiments. To assess error in our posterior model estimates, we used Hamiltonian Monte Carlo (HMC) to perform inference in one of our toy datasets (2D circles). Drawing samples from a 3-component Gaussian Mixture Model (GMM) prior, we integrated the generative process backward in time to obtain corresponding data space samples (A, components shown in orange, blue, and purple). We then used HMC to obtain posterior samples from the posterior distribution over the weights of the GMM components. (B, C) Kernel density estimates from the joint posterior samples over the mixture distribution weights in the dimension-preserving and dimension-reducing cases. Dashed vertical and horizontal lines indicate posterior means for each component. Reference ground-truth weights were $\mathbf{w}=[0.5,0.25,0.25]$.
The only potential source of error, apart from Monte Carlo error, in the above procedure arises from the fact that the score function used in (6) is only an estimate of the true score. To assess whether integrating noisy estimates of the score could produce errant posterior samples, we conducted the experiment showcased in Figure 4A (Appendix B.6). Briefly, we constructed a Gaussian Mixture Model (GMM) prior with three pre-specified components (Appendix B.6) from which we drew samples of $\mathbf{z}$, integrating backwards in time using our trained pfODE networks to construct corresponding observed data points $\mathbf{x}$. We then utilized Hamiltonian Monte Carlo (HMC) 1, 56, 57, 58 to obtain posterior samples for the GMM component weights. As shown in Figure 4B, C, the resulting posterior correctly covers the original ground-truth values, suggesting that numerical errors in score estimates, at least in this simplified scenario, do not appreciably accumulate. This is likely because, empirically, score estimates do not appear to be strongly auto-correlated in time (Appendix C.3), suggesting that $\mathbf{\hat{s}}(\mathbf{x},t)$ is well approximated as a scaled colored noise process and the corresponding pfODE as an SDE. In such a case, standard theorems for SDE integration show that while errors due to noise do accumulate, these can be mitigated by a careful choice of integrator and ultimately controlled by reducing step size [59, 60]. In addition, we verified this empirically in both low-dimensional examples (Figure 4, Appendices B.6, C.2.1) and with round-trip integration of the pfODE in high-dimensional datasets (Tables 1, 2, Appendix B.5.1).
@@ 6 Experiments
For the PR-Reducing flows, the final scale ratio between preserved vs. shrunken dimensions for finite integration times is dependent on the quantity $e^{\rho(g_{*}-g_{i})T}$. Therefore, for fixed end integration time $T$ and rate $\rho$, this scaling is dictated by $g_{*}-g_{i}$, which we call the “inflation gap” (IG), Appendix B.2. As this inflation gap increases, compressed dimensions are shrunken to a greater extent, and the denoising networks are required to amortize score estimation over wider noise scales, a harder learning problem. Therefore, for our proposed model, compression should be understood both in terms of the number of dimensions being preserved and the size of this inflation gap.
To assess how these two factors affect model performance, we performed two sets of experiments on two benchmark image datasets (CIFAR-10 [61] and AFHQv2 [62]; Appendix B.4.2; code: [63]; project website: [64]). In the first set of experiments, we fixed $T$, $\rho$, and the inflation gap ($\mathrm{IG}=1.02$) while varying only the number of preserved dimensions $d$ between $d=1$ (compression to $\approx 0.03\%$) and $d=3072$ (no compression) for both datasets. For the second set of experiments, we worked with the AFHQv2 dataset and fixed $T$, $\rho$, and $d=2$, while varying the inflation gap ($\text{IG}=1.10,1.25,1.35,1.50$). In Tables 1 and 2 we showcase Frechet Inception Distance (FID) scores [65] (mean ±$\pm 2\sigma$ over 3 independently generated sets of images, each with 50,000 samples) and round-trip integration mean squared errors (mean MSE ±$\pm 2\sigma$ over 3 randomly sampled sets of images, each with 10,000 samples) for each ($d$, IG) combination explored (Appendices B.5.1, B.5.2, B.5.4). Figures 5, 6, and 7 showcase 24 randomly generated images (top rows) along with round-trip integration results for 8 randomly sampled images (bottom rows), across select ($d$, IG) combinations.
Finally, we also compared our inflationary flows (IFs) model generative performance on CIFAR-10 against three existing injective flow model baselines (Appendix B.5.2) — M-Flows [21], Rectangular Flows (RFs) [22], and Canonical Manifold Flows (CMF) [23] — for different numbers of preserved dimensions ($d=30,40,62$). Table 3 showcases best FID scores (out of 3 independently generated sets of images, each with 10,000 samples) for each such experiment. For these comparison experiments, we fixed IG=1.02 when training our networks for the different $d$ values.
As a general trend, increasing the number of preserved dimensions at a constant inflation gap led to improvements in generative quality (lower FID scores) and reduced MSE (Table 1). However, some schedules we assessed are not entirely consistent with this trend. We hypothesize this is at least partially due to variance arising from different network initializations for each schedule (Appendix B.5.3), as well as differences between the two datasets explored here. As expected, increasing inflation gap while maintaining the number of preserved dimensions leads to worsened generative performance (higher FID scores, Table 2). Finally, in terms of predictive calibration, our model provides substantial gains when compared to existing injective flow model baselines (Table 3).
@@
Figure 5: Generation and round-trip experiments for AFHQv2 at IG=1.02 and varying number of preserved dimensions. Top row: Generated samples for select flow schedules (PR-Preserving (PRP), PR-Reducing to 2D ($\approx 0.07\%$), 30D($\approx 1\%$), and 307D($\approx 10\%$), at 1.02 IG. Bottom row: Results for round-trip experiments under same schedules. Leftmost columns are original samples, middle columns are samples mapped to Gaussian latent spaces, and rightmost columns are recovered samples.
Table 1: FID and round-trip MSE (mean ±$\pm 2\sigma$) at 1.02 Inflation Gap (IG)
AFHQv2 CIFAR-10
Dimensions FID MSE Dimensions FID MSE
1 $12.65\pm 0.07$ $1.47\pm 0.07$ 1 $20.76\pm 0.09$ $1.07\pm 0.10$
2 $11.95\pm 0.06$ $1.55\pm 0.21$ 2 $21.29\pm 0.04$ $0.81\pm 0.11$
30 $13.64\pm 0.02$ $3.79\pm 0.13$ 30 $23.36\pm 0.14$ $2.21\pm 0.08$
62 $14.05\pm 0.18$ $5.32\pm 0.18$ 62 $23.30\pm 0.19$ $2.27\pm 0.24$
307 $15.64\pm 0.10$ $3.33\pm 0.13$ 307 $28.07\pm 0.13$ $0.71\pm 0.02$
615 $14.63\pm 0.07$ $2.42\pm 0.18$ 615 $24.49\pm 0.27$ $0.29\pm 0.03$
1536 $13.36\pm 0.12$ $0.14\pm 0.03$ 1536 $17.44\pm 0.16$ $0.16\pm 0.06$
3041 $13.97\pm 0.13$ $0.28\pm 0.06$ 3041 $16.60\pm 0.05$ $0.30\pm 0.02$
3072 $11.90\pm 0.08$ $0.38\pm 0.04$ 3072 $17.01\pm 0.10$ $0.22\pm 0.03$
@@
Table 2: FID and round-trip MSE (mean ±$\pm 2\sigma$) for AFHQv2 at varying Inflation Gaps (IG)
Dimensions IG FID MSE
$2$ 1.02 $11.95\pm 0.06$ $1.55\pm 0.21$
$2$ 1.10 $13.98\pm 0.13$ $1.35\pm 0.08$
$2$ 1.25 $17.84\pm 0.15$ $1.65\pm 0.09$
$2$ 1.35 $34.68\pm 0.37$ $1.19\pm 0.18$
$2$ 1.50 $107.64\pm 0.43$ $0.11\pm 0.02$
Table 3: FID score comparison with injective flows for CIFAR-10
Dimensions Preserved IFs (IG=1.02) M-Flow RFs CMFs
$30$ 23.3 541.2 544.0 532.6
$40$ 24.3 535.7 481.3 444.6
$62$ 23.2 280.9 280.8 287.9